# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v1

@@ II System Model And Problem Formulation
${\gamma_{\_}k}=\frac{{{{\bar{P}}_{\_}k}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}k}}\right|}^{2}}}}{{\sum\limits_{\_}{i=1,i\neq k}^{K}{{{\bar{P}}_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}}+1}}=\frac{{{{\bar{P}}_{\_}k}{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}k}{\bf{h}}_{\_}k^{H}{{\bf{v}}_{\_}k}}}{{{\bf{v}}_{\_}k^{H}{{\bf{C}}_{\_}k}{{\bf{v}}_{\_}k}}},$ (8)
where ${{\bf{C}}_{\_}k}\triangleq{{\bf{I}}_{\_}{LM}}+\sum\nolimits_{\_}{i=1,i\neq k}^{K}{{{\bar{P}}_{\_}i}{{\bf{h}}_{\_}i}{\bf{h}}_{\_}i^{H}}$ denotes the interference-plus-noise covariance matrix with respect to (w.r.t.) UE $k$, and ${{\bar{P}}_{\_}i}\triangleq{P_{\_}i}/{\sigma^{2}}$. The achievable rate of UE $k$ is
${R_{\_}k}={\log_{\_}2}\left({1+{\gamma_{\_}k}}\right).$ (9)
Our objective is to maximize the minimum achievable rate over $K$ UEs, by jointly optimizing the 3D UAV swarm placement positions ${\bf{Q}}=\left\{{{{\bf{q}}_{\_}l},\forall l}\right\}$, the local positions of MAs at all UAVs ${\bf{B}}=\left\{{{b_{\_}{l,m}},\forall l,m=2,\cdots,M}\right\}$, and receive beamforming for all UEs ${\bf{V}}=\left\{{{{\bf{v}}_{\_}k},\forall k}\right\}$. This optimization problem can be formulated as
[…]
${\rm{s.t.}}$ $\ {{\bf{q}}_{\_}l}\in{\cal C},\ \forall l,$ (10a)
$\ {\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq d_{\_}{\min}^{2},\ \forall l,l^{\prime}\neq l,$ (10b)
$\ {b_{\_}{l,m}}\in{\tilde{\cal C}},\ \forall l,m,$ (10c)
$\ {\left|{{b_{\_}{l,m}}-{b_{\_}{l,m^{\prime}}}}\right|^{2}}\geq\tilde{d}_{\_}{\min}^{2},\ \forall l,m,m^{\prime}\neq m,$ (10d)
$\ \left\|{{{\bf{v}}_{\_}k}}\right\|=1,\ \forall k,$ (10e)
where $\cal C$ denotes the movable region of UAVs, ${\tilde{\cal C}}$ denotes the (local) movable region of MAs at each UAV, and the collision and mutual coupling avoidance constraints in (1) and (2) have been replaced by their equivalent quadratic forms, i.e., (10b) and (10d), respectively. Problem (P1) is challenging to be directly solved due to the non-concave objective function and (some) non-convex constraints. Besides, the optimization variables ${\bf Q}$, ${\bf B}$ and ${\bf V}$ are intricately coupled in the objective function, which makes their joint optimization a challenging task.
@@ III Special Case of Single-Antenna UAV
$\mathop{\max}\limits_{\_}{{\bf{Q}},{\bf{V}}}$ $\ \mathop{\min}\limits_{\_}{k\in{\cal K}}\ {R_{\_}k}$ (11)
${\rm{s.t.}}$ $\ {{\bf{q}}_{\_}l}\in{\cal C},\ \forall l,$
$\ {\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq d_{\_}{\min}^{2},\ \forall l,l^{\prime}\neq l,$
$\ \left\|{{{\bf{v}}_{\_}k}}\right\|=1,\ \forall k.$
@@ III-A Single UE and Two UEs
which only depends on the number of UAVs $L$, irrespective of the geometry formed by the UAV swarm enabled MA system. As a result, any UAV swarm placement satisfying the collision avoidance constraint can achieve the maximum achievable rate.
Next, we consider the classic MRC, zero-forcing (ZF) and minimum mean-square error (MMSE) beamforming schemes for the case of two UEs. Denote by ${\xi_{\_}{k,k^{\prime}}}\triangleq\frac{{{{\left|{{\bf{h}}_{\_}k^{H}{{\bf{h}}_{\_}{k^{\prime}}}}\right|}^{2}}}}{{{{\left\|{{{\bf{h}}_{\_}k}}\right\|}^{2}}{{\left\|{{{\bf{h}}_{\_}{k^{\prime}}}}\right\|}^{2}}}}$ the channel’s squared-correlation coefficient between UEs $k$ and $k^{\prime}$, where $k,k^{\prime}=1,2$ and $k\neq k^{\prime}$. Under the far-field UPW model, we have ${\left\|{{{\bf{h}}_{\_}k}}\right\|^{2}}={\left|{{\alpha_{\_}k}}\right|^{2}}L$. The resulting SINR/SNR of UE $k$ after applying the three beamforming schemes are respectively given by [34, 35]
${\gamma_{\_}{{\rm{MRC}},k}}={{\bar{P}}_{\_}k}{\left|{{\alpha_{\_}k}}\right|^{2}}L\left({1-\frac{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L{\xi_{\_}{k,k^{\prime}}}}}{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L{\xi_{\_}{k,k^{\prime}}}+1}}}\right),$ (14)
${\gamma_{\_}{{\rm{ZF}},k}}={{\bar{P}}_{\_}k}{\left|{{\alpha_{\_}k}}\right|^{2}}L\left({1-{\xi_{\_}{k,k^{\prime}}}}\right),$ (15)
${\gamma_{\_}{{\rm{MMSE}},k}}={{\bar{P}}_{\_}k}{\left|{{\alpha_{\_}k}}\right|^{2}}L\left({1-\frac{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L{\xi_{\_}{k,k^{\prime}}}}}{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L+1}}}\right).$ (16)
It is observed from (14)-(16) that the SINRs of MRC, ZF, and MMSE beamforming depend on the coefficient ${\xi_{\_}{k,k^{\prime}}}$, and a larger SINR can be achieved as ${\xi_{\_}{k,k^{\prime}}}$ decreases. Moreover, the terms ${\frac{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L{\xi_{\_}{k,k^{\prime}}}}}{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L{\xi_{\_}{k,k^{\prime}}}+1}}}$, ${\xi_{\_}{k,k^{\prime}}}$, and ${\frac{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L{\xi_{\_}{k,k^{\prime}}}}}{{{{\bar{P}}_{\_}{k^{\prime}}}{{\left|{{\alpha_{\_}{k^{\prime}}}}\right|}^{2}}L+1}}}$ in (14)-(16) account for the SNR loss factors for UE $k$ due to applying the MRC, ZF, and MMSE beamforming schemes, respectively. By substituting (5) into ${\xi_{\_}{k,k^{\prime}}}$, we have
${\xi_{\_}{k,k^{\prime}}}=\frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{\_}k^{H}{{\bf{a}}_{\_}{k^{\prime}}}}\right|^{2}}.$ (17)
@@
$\mathop{\min}\limits_{\_}{\bf{Q}}$ $\ \frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{\_}k^{H}{{\bf{a}}_{\_}{k^{\prime}}}}\right|^{2}}$ (18)
${\rm{s.t.}}$ $\ {{\bf{q}}_{\_}l}\in{\cal C},\ \forall l,$
$\ {\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq d_{\_}{\min}^{2},\ \forall l,l^{\prime}\neq l.$
To tackle problem (18), we first study the property of the objective function. By substituting (6) into the objective function, we have
$\frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{\_}k^{H}{{\bf{a}}_{\_}{k^{\prime}}}}\right|^{2}}$ $=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{\_}{l=1}^{L}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\varphi_{\_}k}\left({{{\bf{q}}_{\_}l}}\right)-{\varphi_{\_}{k^{\prime}}}\left({{{\bf{q}}_{\_}l}}\right)}\right)}}}}\right|^{2}}$ (19)
$=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{\_}{l=1}^{L}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{{\left({{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right)}^{T}}{{\bf{q}}_{\_}l}}}}}\right|^{2}}.$
By letting ${\nu_{\_}{\min}}={\left\lceil{L{d_{\_}{\min}}\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|/\lambda-1}\right\rceil}$, we obtain the following theorem.
@@
${{\bf{q}}_{\_}l}={{\bf{q}}_{\_}1}+\left({l-1}\right)\frac{{\left({{\varsigma^{\star}}+1/L}\right)\lambda}}{{{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}^{2}}}}\left({{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right),$ (20)
where ${{\bf{q}}_{\_}1}$ can be any vector to guarantee ${{\bf{q}}_{\_}l}\in{\cal C}$, $\forall l$, and ${\varsigma^{\star}}$ is given by
${\varsigma^{\star}}=\left\{\begin{split}&\frac{{{\nu_{\_}{\min}}}}{L},\ \ \ \ \ \ {\rm if}\ \bmod\left({{\nu_{\_}{\min}}+1,L}\right)\neq 0,\\ &\frac{{{\nu_{\_}{\min}}+1}}{L},\ {\rm otherwise}.\end{split}\right.$ (21)
@@
Note that Theorem 1 extends the 1D result in [23] to the general 3D placement, and the derived closed-form solution ${\varsigma^{\star}}$ can be a fraction, which removes the requirement of being an integer imposed in [23]. Moreover, UAV placement position ${{\bf{q}}_{\_}l}\in{\cal C}$, $\forall l$, can be satisfied when $\left\|{{{\bf{q}}_{\_}L}-{{\bf{q}}_{\_}1}}\right\|=\frac{{\left({L-1}\right)\left({{\varsigma^{\star}}+1/L}\right)\lambda}}{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}}$ is no greater than the maximum distance between any two points in ${\cal C}$.
With (20) and (21), we have ${\xi_{\_}{k,k^{\prime}}}=0$, and ${\gamma_{\_}{{\rm{MRC}},k}}={\gamma_{\_}{{\rm{ZF}},k}}\ ={\gamma_{\_}{{\rm{MMSE}},k}}={{\bar{P}}_{\_}k}{\left\|{{{\bf{h}}_{\_}k}}\right\|^{2}}={{\bar{P}}_{\_}k}{\left|{{\alpha_{\_}k}}\right|^{2}}L$, i.e., they are identical to the single UE SNR without IUI. Thus, an IUI-free communication can be achieved with the UAV swarm enabled MA system, by setting the UAVs’ static positions according to Theorem 1. It is worth noting that the inter-UAV distance is $\frac{{\left|{{\varsigma^{\star}}+1/L}\right|\lambda}}{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}}\geq d_{\_}{\min}$, where the minimum distance $d_{\_}{\min}$ is in general much larger than half wavelength. Thus, the UAV swarm forms a USA for the IUI-free communication. Moreover, the objective function in (P1) is given by
$\mathop{\min}\limits_{\_}{k\in K}{\log_{\_}2}\left({1+{{\bar{P}}_{\_}k}{{\left|{{\alpha_{\_}k}}\right|}^{2}}L}\right).$ (22)

### v2

@@ II System Model And Problem Formulation
${\gamma_{k}}=\frac{{{{\bar{P}}_{k}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{k}}}\right|}^{2}}}}{{\sum\limits_{i=1,i\neq k}^{K}{{{\bar{P}}_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}}+1}}=\frac{{{{\bar{P}}_{k}}{\bf{v}}_{k}^{H}{{\bf{h}}_{k}}{\bf{h}}_{k}^{H}{{\bf{v}}_{k}}}}{{{\bf{v}}_{k}^{H}{{\bf{C}}_{k}}{{\bf{v}}_{k}}}},$ (9)
where ${{\bf{C}}_{k}}\triangleq{{\bf{I}}_{LM}}+\sum\nolimits_{i=1,i\neq k}^{K}{{{\bar{P}}_{i}}{{\bf{h}}_{i}}{\bf{h}}_{i}^{H}}$ denotes the interference-plus-noise covariance matrix with respect to (w.r.t.) UE $k$, and ${{\bar{P}}_{i}}\triangleq{P_{i}}/{\sigma^{2}}$. The achievable rate of UE $k$ is
${R_{k}}={\log_{2}}\left({1+{\gamma_{k}}}\right).$ (10)
Our objective is to maximize the minimum achievable rate over $K$ UEs, by jointly optimizing the 3D UAV swarm (relative) placement positions ${\bf{Q}}=\left\{{{{\bf{q}}_{l}},\forall l}\right\}$, the local positions of MAs at all UAVs ${\bf{B}}=\left\{{{b_{l,m}},\forall l,m=2,\cdots,M}\right\}$, and receive beamforming for all UEs ${\bf{V}}=\left\{{{{\bf{v}}_{k}},\forall k}\right\}$. Let $\rho\left({{\bf{Q}},{\bf{B}},{\bf{V}}}\right)=\mathop{\min}\limits_{k\in{\cal K}}{R_{k}}$, which is a function of $\bf{Q}$}, $\bf{B}$}, and $\bf{V}$}. The optimization problem can be formulated as
$\left({\rm{P1}}\right)\ \$$\mathop{\max}\limits_{{\bf{Q}},{\bf{B}},{\bf{V}},\rho}\ \rho$ (11)
${\rm{s.t.}}$ $\ {R_{k}}\geq\rho,\ \forall k,$ (11a)
$\ {{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C},\ \forall l,$ (11b)
$\ {\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}\geq d_{\min}^{2},\ \forall l,l^{\prime}\neq l,$ (11c)
$\ {b_{l,m}}\in{\tilde{\cal C}},\ \forall l,m,$ (11d)
$\ {\left|{{b_{l,m}}-{b_{l,m^{\prime}}}}\right|^{2}}\geq\tilde{d}_{\min}^{2},\ \forall l,m,m^{\prime}\neq m,$ (11e)
$\ \left\|{{{\bf{v}}_{k}}}\right\|=1,\ \forall k,$ (11f)
where $\cal C$ denotes the movable region of UAVs, $\tilde{\cal C}$} denotes the (local) movable region of MAs at each UAV, and the minimum safe distance constraint and mutual coupling avoidance constraint in (1) and (2) have been replaced by their equivalent quadratic forms, i.e., (11c) and (11e), respectively. Problem (P1) is challenging to be directly solved due to the non-concave objective function and (some) non-convex constraints. Besides, the optimization variables $\bf Q$}, $\bf B$} and $\bf V$} are intricately coupled in the objective function, which makes their joint optimization a challenging task.
@@ III Special Case of Single-Antenna UAV
$\mathop{\max}\limits_{{\bf{Q}},{\bf{V}},\rho}$ $\ \rho$ (12)
${\rm{s.t.}}$ $\ {R_{k}}\geq\rho,\ \forall k,$
$\ {{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C},\ \forall l,$
$\ {\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}\geq d_{\min}^{2},\ \forall l,l^{\prime}\neq l,$
$\ \left\|{{{\bf{v}}_{k}}}\right\|=1,\ \forall k.$
@@ III-A Single UE and Two UEs
which only depends on the number of UAVs $L$, irrespective of the geometry formed by the UAV swarm enabled MA system. As a result, any UAV swarm placement within the movable region while satisfying the minimum safe distance constraint can achieve the maximum achievable rate.
Next, we consider the classic MRC, zero-forcing (ZF) and minimum mean-square error (MMSE) beamforming schemes for the case of two UEs. Denote by ${\xi_{k,k^{\prime}}}\triangleq\frac{{{{\left|{{\bf{h}}_{k}^{H}{{\bf{h}}_{k^{\prime}}}}\right|}^{2}}}}{{{{\left\|{{{\bf{h}}_{k}}}\right\|}^{2}}{{\left\|{{{\bf{h}}_{k^{\prime}}}}\right\|}^{2}}}}$ the channel’s squared-correlation coefficient between UEs $k$ and $k^{\prime}$, where $k,k^{\prime}=1,2$ and $k\neq k^{\prime}$. Under the far-field UPW model, we have ${\left\|{{{\bf{h}}_{k}}}\right\|^{2}}={\left|{{\alpha_{k}}}\right|^{2}}L$. The resulting SINR/SNR of UE $k$ after applying the three beamforming schemes are respectively given by [34, 35]
${\gamma_{{\rm{MRC}},k}}={{\bar{P}}_{k}}{\left|{{\alpha_{k}}}\right|^{2}}L\left({1-\frac{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L{\xi_{k,k^{\prime}}}}}{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L{\xi_{k,k^{\prime}}}+1}}}\right),$ (15)
${\gamma_{{\rm{ZF}},k}}={{\bar{P}}_{k}}{\left|{{\alpha_{k}}}\right|^{2}}L\left({1-{\xi_{k,k^{\prime}}}}\right),$ (16)
${\gamma_{{\rm{MMSE}},k}}={{\bar{P}}_{k}}{\left|{{\alpha_{k}}}\right|^{2}}L\left({1-\frac{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L{\xi_{k,k^{\prime}}}}}{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L+1}}}\right).$ (17)
It is observed from (15)-(17) that the SINRs of MRC, ZF, and MMSE beamforming depend on the coefficient $\xi_{k,k^{\prime}}$}, and a larger SINR can be achieved as $\xi_{k,k^{\prime}}$} decreases. Moreover, the terms $\frac{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L{\xi_{k,k^{\prime}}}}}{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L{\xi_{k,k^{\prime}}}+1}}$}, $\xi_{k,k^{\prime}}$}, and $\frac{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L{\xi_{k,k^{\prime}}}}}{{{{\bar{P}}_{k^{\prime}}}{{\left|{{\alpha_{k^{\prime}}}}\right|}^{2}}L+1}}$} in (15)-(17) account for the SNR loss factors for UE $k$ due to applying the MRC, ZF, and MMSE beamforming schemes, respectively. By substituting (6) into $\xi_{k,k^{\prime}}$}, we have
${\xi_{k,k^{\prime}}}=\frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{k}^{H}{{\bf{a}}_{k^{\prime}}}}\right|^{2}}.$ (18)
@@
$\mathop{\min}\limits_{\bf{Q}}$ $\ \frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{k}^{H}{{\bf{a}}_{k^{\prime}}}}\right|^{2}}$ (19)
${\rm{s.t.}}$ $\ {{\bf{q}}_{l}}\in{\cal C},\ \forall l,$
$\ {\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}\geq d_{\min}^{2},\ \forall l,l^{\prime}\neq l.$
To tackle problem (19), we first study the property of the objective function. Note that when the movable region size is much smaller than the link distance, the UPW model gives a valid approximation of spherical wavefront for channels between UAVs and ground UEs [35], which may correspond to the cases of small movable region and/or high UAV operation altitude. In this case, the receive response vector in (7) reduces to ${{\bf{a}}_{k}}={e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{r_{k}}}}{[{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{\bf{q}}_{1}}}\right)}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{\bf{q}}_{L}}}\right)}}}]^{T}}$, where ${\varphi_{k}}\left({{{\bf{q}}_{l}}}\right)={\bm{\kappa}}_{k}^{T}{{\bf{q}}_{l}}$. By substituting it into the objective function, we have
$\frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{k}^{H}{{\bf{a}}_{k^{\prime}}}}\right|^{2}}$ $=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\varphi_{k}}\left({{{\bf{q}}_{l}}}\right)-{\varphi_{k^{\prime}}}\left({{{\bf{q}}_{l}}}\right)}\right)}}}}\right|^{2}}$ (20)
$=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{{\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right)}^{T}}{{\bf{q}}_{l}}}}}}\right|^{2}}.$
By letting ${\nu_{\min}}={\left\lceil{L{d_{\min}}\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|/\lambda-1}\right\rceil}$, we obtain the following theorem for UAV position without drift. In particular, high-accuracy positioning technologies, such as real-time kinematic (RTK) global navigation satellite system (GNSS), multi-sensor fusion, and robust flight control algorithms help to mitigate the issue of UAV position drift [46].
@@
${{\bf{q}}_{l}}={{\bf{q}}_{1}}+\left({l-1}\right)\frac{{\left({{\varsigma^{\star}}+1/L}\right)\lambda}}{{{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}^{2}}}}\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right),$ (21)
where ${\bf{q}}_{1}$} can be any vector to guarantee ${{\bf{q}}_{l}}\in{\cal C}$, $\forall l$, and $\varsigma^{\star}$} is given by
${\varsigma^{\star}}=\left\{\begin{split}&\frac{{{\nu_{\min}}}}{L},\ \ \ \ \ \ {\rm if}\ \bmod\left({{\nu_{\min}}+1,L}\right)\neq 0,\\ &\frac{{{\nu_{\min}}+1}}{L},\ {\rm otherwise}.\end{split}\right$. (22)
@@
Note that Theorem 1 extends the 1D result in [23] to the general 3D placement, and the derived closed-form solution $\varsigma^{\star}$} can be a fraction, which removes the requirement of being an integer imposed in [23]. Moreover, UAV placement position ${{\bf{q}}_{l}}\in{\cal C}$, $\forall l$, can be satisfied when $\left\|{{{\bf{q}}_{L}}-{{\bf{q}}_{1}}}\right\|=\frac{{\left({L-1}\right)\left({{\varsigma^{\star}}+1/L}\right)\lambda}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}$ is no greater than the maximum distance between any two points in $\cal C$}.
With (21) and (22), we have ${\xi_{k,k^{\prime}}}=0$, and ${\gamma_{{\rm{MRC}},k}}={\gamma_{{\rm{ZF}},k}}\ ={\gamma_{{\rm{MMSE}},k}}={{\bar{P}}_{k}}{\left\|{{{\bf{h}}_{k}}}\right\|^{2}}={{\bar{P}}_{k}}{\left|{{\alpha_{k}}}\right|^{2}}L$, i.e., they are identical to the single UE SNR without IUI. Thus, an IUI-free communication can be achieved with the UAV swarm enabled MA system, by setting the UAVs’ static positions according to Theorem 1. It is worth noting that the inter-UAV distance is $\frac{{\left|{{\varsigma^{\star}}+1/L}\right|\lambda}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}\geq d_{\min}$, where the minimum distance $d_{\min}$ is in general much larger than half wavelength. Thus, the UAV swarm forms a USA for the IUI-free communication. In particular, compared to the classic compact array where antenna spacing is separated by half wavelength, USA can achieve a larger array aperture, and thus a narrower beamwidth of the main lobe is enabled.
Moreover, the objective function in (P1) is given by
$\mathop{\min}\limits_{k\in{\cal K}}{\log_{2}}\left({1+{{\bar{P}}_{k}}{{\left|{{\alpha_{k}}}\right|}^{2}}L}\right).$ (23)